Schedule September 2008

return to main 1401 Restoration Page
go to Team Bios

Contents:

Ed took a vacation from posting weekly details -
Wed September 24 - general, Thurs September 25 - tape team


Wed September 24 - general

Present were: Ron Williams, Bob Erickson, Frank King, Allen Palmer, Glenn Lea, Joe Preston, Stan Paddock, Judith Haemmerle, Ed Thelen, Robert Garner

Following report from Robert Garner


Status

  1. DE 1401
    • Fully functional and 100% operational, for several consecutive months now! (Not to be upstaged by the new-kid CT 1401. ;-)
    • Allen noted defective power supply on 2nd Model V 729. 3V instead of 12V. Ed investigating.
  2. CT 1401



Thurs September 25 - tape team

From Bob Feretich
1401 TAU debug status 9/25/08 (Jeff & I)

We replaced 2 burnt out indicator lamps and 3 bad cards.

  • We are now able to write records to tape from the TAU CE panel. The data was received correctly by the PC. The TAU was happy with the wrap read back of the write data. No errors are reported.
  • The Rewind switch also seems to work.
  • The Rewind & Unload switch does not work.
Regards,
Bob


Sat. September 27 - general

Attendees: Ron Williams, Bob Erickson
  1. Worked on 1402 read check problem in 1401 logic

  2. Bob noted that the 1402, after decades of use, needs a thorough cleaning and rehab. (As soon as CT system is up, and since the DE 1401 is operational, we should consider a full mechanical overhaul/refurb of the CT 1402!)



Go to
October 2008
go to Team Bios